OpenStudy (anonymous):

Will fan and give medals What are the zero(s) of the function f(x)= 5x^2-25/x?

OpenStudy (anonymous):

\[f(x)=\frac{5x^2-25}{x}\] right?

OpenStudy (anonymous):

Yes(: @satellite73

OpenStudy (anonymous):

a fraction is only zero if the numerator is

OpenStudy (anonymous):

set \[5x^2-25=0\] and solve for \(x\) in three steps

OpenStudy (anonymous):

\[5x^2-25=0\\ 5x^2=25\\ x^2=5\\ x=\pm\sqrt5\]

OpenStudy (anonymous):

 If the numerator becomes 0, the fraction will turn to 0 , If the denominator become 0, the result is undefined.